Solve the system using elimination. Make sure to show all your work for full credit.

5x + 4y = 1
3x - 4y = 7

Answer by amarjeeth123(569)

The given equations are:
5x + 4y = 1........equation 1
3x - 4y = 7........equation 2
Adding 1 and 2 we get,
5x+3x=1+7
8x=8
x=1.
y=(1-5)/4=-1.
The solution is (1,-1).